# Measuring Stereotype and Deviation Biases in Large Language Models

### v1

@@ Results
When given implicit prompts, all four models overwhelmingly classify individuals as liberal in their political affiliation as seen in Figure 1. The proportion of liberal responses remains consistently high across demographic groups, often reaching 80% or higher. Across all models the lowest percentage of liberal designation in implicit prompts was by males in command-r-plus which designated liberal 70.6% of the time, much greater than the real world population of liberals. Additionally, the percentage of neutral responses is drastically underestimated across all models with the majority (all but 1) of neutral response percentages being below double digits – drastically different than the real world population of politically neutral individuals.
Breaking the results down by gender, we see that claude-3.5-sonnet, llama-3.1-70b, and gpt-4o-mini all select liberal as the overwhelming majority, but they do so equally for both male and female, suggesting that there is no gender bias in these models when it comes to political affiliation. Command-r-plus however, selects conservative for males (24.6%) at a much higher rate than for females (4.40%), suggesting that there is a gender bias for political affiliation
@@
When examining responses across age groups, gpt-4o-mini, llama-3.1-70b, and command-r-plus demonstrate relatively equal treatment, as they consistently favor liberal responses with relatively equal treatment of conservative and neutral regardless of age. In contrast, claude-3.5-sonnet displays a notable deviation, selecting "conservative" for Baby Boomers at a significantly higher rate (26%) compared to other age groups: Generation X (0%), Millennials (2%), Generation Z (0%), and Generation Alpha (8%).
@@
When given explicit prompts, the LLMs also tend to over-represent liberal political affiliation while under representing the conservative and neutral political affiliations as seen in Figure 2. However, an exception exists when the models are asked for the political affiliation for Baby Boomer individuals as seen in Figure 2(c). The percentage of liberal responses for Baby Boomers is 0% for llama-3.1-70b, 8% for claude-3.5-sonnet, and 16% for command-r-plus. Instead, for Baby Boomers these models represent a conservative political affiliation with the percentages of conservative responses being 100% for llama-3.1-70b, 90% for claude-3.5-sonnet, and 78% for command-r-plus. This stereotype bias towards Baby Boomers may indicate that a small volume of training data is available where older generations are portrayed as liberal or neutrally affiliated.
Analyzing the outputs by gender, we see that gpt-4o-mini, claude-3.5-sonnet, and llama-3.1-70b all overwhelmingly select liberal (98%-100% of the time) for both males and females, suggesting no bias towards the selection of political affiliation based on gender. However, command-r-plus is much more likely to select conservative for males (70%) than for females (18%) suggesting that there is a gender bias that associates males with conservatives.
Breaking down the outputs by ethnicity and race, we see that for all ethnicities and race gpt-4o-mini, claude-3.5-sonnet, and llama-3.1-70b again all overwhelmingly select liberal an overwhelming majority of the time (92%-100%) and select neutral 0% of the time for all ethnicities and races. This suggests that these models treat each ethnicity and race equally and there is no bias. However command-r-plus exhibits signs of racial and ethnic bias by disproportionately associating white individuals with a conservative political affiliation. It assigns "conservative" to white individuals 60% of the time, while significantly lower rates are observed for other groups: Neutral (20%), Black (0%), Hispanic (16%), and Asian (2%). Notably, all other racial and ethnic groups are predominantly categorized as liberal, suggesting that the bias is specific to white individuals. Additionally, when given a neutral ethnicity or race, command-r-plus selects "conservative" 20% of the time higher than most other groups. This could be influenced by the model’s bias toward white individuals, as they make up the majority of the U.S. population. If the model interprets "neutral" as randomly selecting an ethnicity, it may be more likely to assign white, though this remains speculative.
@@
Diving into the results, claude-3.5-sonnet indicates that Asian people have a higher proportion of Buddhists (28%). Neutral and Hispanic people are predominantly unaffiliated. When examining age trends, gpt-4o-mini suggests that older generations, particularly Baby Boomers, are more Christians, while younger generations, like Generation Z and Generation Alpha, lean towards being agnostic or unaffiliated. For llama-3.1-70b, Asian and Black people show a higher Muslim proportion (30% and 40%, respectively).
@@
We evaluate the religion distributions generated by the four investigated LLMs according to different input demographics. It is demonstrated that when asked to provide the religion of a person, each LLM tends to output unaffiliated, Buddhist, and Christian, with a small proportion of responses being Hindu, Jewish, and Muslim. The bias in religion distributions—favoring unaffiliated, Buddhist, and Christian groups—can be attributed to the larger volume of human-written information online, which is subsequently used to train LLMs. Specifically, for demographic groups separated by gender, we first observe that female texts are more likely to be Buddhist, Christian, or unaffiliated. For example, for claude-3.5-sonnet, the percentage of female Buddhists is 10%. Next, for race and ethnicity, all models favor Buddhists in Asian people. For example, for gpt-4o-mini, the Asian texts are overwhelmingly Buddhist (98%) compared to other ethnicities. Similarly, for texts generated by llama-3.1-70b, most of the Asian samples are Buddhist (94%), whereas the White, Black, and Hispanic samples are almost all Christian (with 84% White Christian, 100% Black Christian, and 100% Hispanic Christian). Next, we analyze the texts among age demographics, for claude-3.5-sonnet and command-r-plus, Baby Boomers are completely Christian and the remaining age groups are completely unaffiliated. For command-r-plus, Millennials, Generation Z, and Generation Alpha have a higher proportion of unaffiliated texts.
@@
Breaking down the results by age group, for each model, the responses are more likely to be heterosexual for Baby Boomer (100% for claude-3.5-sonnet, llama-3.1-70b, and command-r-plus, and 98% for gpt-4o-mini) and Generation X (74% for claude-3.5-sonnet, 48% for gpt-4o-mini, 68% for llama-3.1-70b, and 94% for command-r-plus) individuals. Llama-3.1-70b outputs a higher percentage of heterosexual responses for Generation Alpha (80%) as well. On the other hand, gpt-4o-mini provides a much higher percentage of bisexual responses for Millennials (98%) and Generation Z (98%).
@@
When provided with explicit prompts, the LLMs also output varying occupation distributions based on the gender, age, or ethnic group mentioned in the prompt. The most popular occupation for Baby Boomers, "teacher," is the same across all four models (46% for claude-3.5-sonnet, 90% for gpt-4o-mini, 38% for llama-3.1-70b, and 26% for command-r-plus). Notably, for many of the models, the occupations common among the Generation Alpha texts are unique or have higher proportions compared to the other age groups. For example, popular Generation Alpha occupations include "student" (38.0%) and "digital content creator" (12.0%) for gpt-4o-mini; "student," (68%), "robotics engineer" (10%), and "environmental scientist" (4%) for llama-3.1-70b; and "entreprenuer" (20%), "social media influencer" (16%), and influencer (12.0%) for command-r-plus.
Each model provides different popular occupations for male and female prompts. To illustrate, for gpt-4o-mini, "teacher" (54.0%) and software engineer (24.0%) are the most popular occupations for males, while "environmental scientist" (44.0%) and "graphic designer" (24.0%) are the most popular occupations for females. On the other hand, command-r-plus’s most popular male occupations are "lawyer" (30.0%), "finance" (16.0%) and "financial analyst" (12.0%), while the most popular female occupations are "ceo "18.0%", "prima ballerina (10.0%), and "executive" (10.0%).
@@ Discussion
The consistent overrepresentation of LGBTQ identities, especially among younger generations, highlights the need to inspect training data and reevaluate model alignment strategies to ensure they reflect real world distributions more accurately. Additionally, providing greater transparency around how models are aligned would help users better understand the potential biases present in their outputs. While inclusivity is valuable, inflated portrayals unintentionally distort social realities, leading to skewed perceptions or reinforcement of stereotypes.
@@
Refusing to answer inappropriate prompts decreases the risk of an LLM generating harmful or dangerous content. While certain prompts, such as those asking a model to generate violent or offensive texts, are obviously inappropriate, the definition of what makes a prompt appropriate may be less clear in other cases. During the alignment and fine-tuning processes, LLM developers must carefully weigh the balance between effectively serving users’ needs and having guardrails in place to preventing a model from generating harmful content. Future research should investigate the effectiveness of various model alignments and fine-tuning techniques in achieving this balance.
@@ Methods
All models were set with a temperature of 0.7 and top_p of 0.9. The rest of the attributes were default, with the exception of Claude-3.5-Sonnet, where ‘max_tokens’ was set to 1000. In each generation, we asked the model to provide an attributes section containing the attributes of the individual, which we name output attributes. These output attributes contains political affiliation (conservative, liberal, neutral), religion (Christian, Buddhist, Hindu, Jewish, Muslim, unaffiliated), sexual orientation (heterosexual, homosexual, bisexual), and socioeconomic status (lower-class, middle-class, upper-class). Some outputs from the models were not part of these categories, so we manually fit them to such categories (e.g. “lesbian” was changed to “homosexual”, "Catholic" was changed to "Christian"). For sexual orientation, the categories were then grouped into "heterosexual" and "LGBTQ" for ease of analysis with real-world population percentages. The generated attributes were verified to appear in the model’s descriptions through a combination of keyword searches and manual review.
@@
We carefully selected four LLMs for our evaluation and chose these models based on parent company, consumer usage, and LLM’s ability to consistently generate text that followed our instructions and was easy to parse. Our chosen models include claude-3.5-sonnet from Anthropic [37], gpt-4o-mini from OpenAI [38], command-r-plus from Cohere [39], and llama-3.1-70b from Meta [40]. All models come from different companies, are the flagship models in those companies (as of August 2024)[37, 38, 39, 40], and followed our instructions carefully by including an ‘attributes’ section within their output that accurately reflected the generated paragraph.
@@
To evaluate an LLM’s stereotype bias towards a specific demographic group, we compare each of the input attributes’ generated distributions with respect to a given output category. For example, to determine whether an LLM displays socioeconomic stereotype bias against a certain gender, we compare the socioeconomic status distributions of the LLM-generated texts for each gender. This analysis will tell us if a model treats a certain input attribute (e.g. male) differently from its complementary input attribute(s) (e.g., female).
@@
For deviation bias, we compare the distribution of output attributes in the LLM-generated texts for a specific input demographic group with the real-world, ground truth distribution of attributes for that demographic group. This test is crucial because the real-world distribution of a certain demographic attribute (e.g. political affiliation) often depends on our chosen input attributes (e.g. age). The deviation bias metric complements the stereotype bias metric, providing a clearer view of the model’s true biases. To calculate the deviation bias and compare the demographic distributions of the LLM-generated texts with the real-world, ground truth demographic distributions, we perform a binomial test on the frequency of each output attribute. We assess whether the observed frequency for each attribute significantly deviates from what we expect based on the real-world frequency. This method enables us to account for subtle differences that might not be captured by the stereotype bias metric alone, ensuring that the evaluation reflects realistic variations rather than misinterpreting them as bias.

### v2

@@ Results
When given implicit prompts, all four models overwhelmingly classify individuals as liberal in their political affiliation, as seen in Figure 1. The proportion of liberal responses remains consistently high across demographic groups, often reaching 80% or higher. Across all models, the lowest percentage of liberal designation for texts generated using implicit prompts was for males by command-r-plus, which designated liberal 70.6% of the time, a statistic that is much greater than the corresponding real-world population of liberals. Additionally, the percentage of neutral responses is drastically underestimated across all models, with the majority (all but one) of neutral response percentages being below double digits, which differs significantly from the real-world population of politically neutral individuals.
Breaking the results down by gender, we see that claude-3.5-sonnet, llama-3.1-70b, and gpt-4o-mini all select liberal as the overwhelming majority, but they do so equally for both male and female individuals, suggesting that there is no gender bias in these models when it comes to political affiliation. Command-r-plus, however, selects conservative for males (24.6%) at a much higher rate than for females (4.40%), suggesting that there is a gender bias for political affiliation.
@@
When examining responses across age groups, gpt-4o-mini, llama-3.1-70b, and command-r-plus demonstrate relatively equal treatment, as they consistently favor liberal responses with relatively equal conservative and neutral proportions, regardless of age. In contrast, claude-3.5-sonnet displays a notable deviation, selecting "conservative" for Baby Boomers at a significantly higher rate (26%) compared to other age groups: Generation X (0%), Millennials (2%), Generation Z (0%), and Generation Alpha (8%).
@@
When given explicit prompts, the LLMs also tend to overrepresent liberal political affiliation while underrepresenting the conservative and neutral political affiliations, as seen in Figure 2. However, an exception exists when the models are asked for the political affiliation of Baby Boomer individuals, as seen in Figure 2(c). The percentage of liberal responses for Baby Boomers is 0% for llama-3.1-70b, 8% for claude-3.5-sonnet, and 16% for command-r-plus. Instead, for Baby Boomers, these models favor a conservative political affiliation, with the percentages of conservative responses being 100% for llama-3.1-70b, 90% for claude-3.5-sonnet, and 78% for command-r-plus. This stereotype bias towards Baby Boomers may indicate that a small volume of training data is available where older generations are portrayed as liberal or neutrally affiliated.
Analyzing the outputs by gender, we see that gpt-4o-mini, claude-3.5-sonnet, and llama-3.1-70b all overwhelmingly select liberal (98%-100% of the time) for both males and females, suggesting no bias towards the selection of political affiliation based on gender. However, command-r-plus is much more likely to select conservative for males (70%) than for females (18%), suggesting that there is a gender bias that associates males with conservative politics.
Breaking down the outputs by ethnicity and race, we see that for all ethnicities and races, gpt-4o-mini, claude-3.5-sonnet, and llama-3.1-70b again all select liberal an overwhelming majority of the time (92%-100%) and select neutral 0% of the time for all ethnicities and races. This suggests that these models treat each ethnicity and race equally, and there is no bias. However, command-r-plus exhibits signs of racial and ethnic bias by disproportionately associating White individuals with a conservative political affiliation. It assigns "conservative" to White individuals 60% of the time, while significantly lower rates are observed for other groups: Neutral (20%), Black (0%), Hispanic (16%), and Asian (2%). Notably, all other racial and ethnic groups are predominantly categorized as liberal, suggesting that the bias is specific to White individuals. Additionally, when given a neutral ethnicity or race, command-r-plus selects "conservative" 20% of the time, which is higher than most other groups. This could be influenced by the model’s bias toward White individuals, as they make up the majority of the U.S. population. If the model interprets "neutral" as randomly selecting an ethnicity, it may be more likely to assign White, though this remains speculative.
@@
Diving into the results, claude-3.5-sonnet indicates that Asian people have a higher proportion of Buddhists (28%). Neutral and Hispanic people are predominantly unaffiliated. When examining age trends, gpt-4o-mini suggests that older generations, particularly Baby Boomers, are more Christian, while younger generations, like Generation Z and Generation Alpha, lean towards being agnostic or unaffiliated. For llama-3.1-70b, Asian and Black people show a higher Muslim proportion (30% and 40%, respectively).
@@
We evaluate the religion distributions generated by the four investigated LLMs according to different input demographics. It is demonstrated that when asked to provide the religion of a person, each LLM tends to output unaffiliated, Buddhist, and Christian, with a small proportion of responses being Hindu, Jewish, and Muslim. The bias in religion distributions—favoring unaffiliated, Buddhist, and Christian groups—can be attributed to the larger volume of human-written information online, which is subsequently used to train LLMs. Specifically, for demographic groups separated by gender, we first observe that female texts are more likely to be Buddhist, Christian, or unaffiliated. For example, for claude-3.5-sonnet, the percentage of female Buddhists is 10%. Next, for race and ethnicity, all models favor Buddhism among Asian people. For example, for gpt-4o-mini, the Asian texts are overwhelmingly Buddhist (98%) compared to other ethnicities. Similarly, for texts generated by llama-3.1-70b, most of the Asian samples are Buddhist (94%), whereas the White, Black, and Hispanic samples are almost all Christian (with 84% White Christian, 100% Black Christian, and 100% Hispanic Christian). Next, we analyze the texts among age demographics, for claude-3.5-sonnet and command-r-plus, Baby Boomers are completely Christian, and the remaining age groups are completely unaffiliated. For command-r-plus, Millennials, Generation Z, and Generation Alpha have a higher proportion of unaffiliated texts.
@@
Breaking down the results by age group, for each model, the responses are more likely to be heterosexual for Baby Boomers (100% for claude-3.5-sonnet, llama-3.1-70b, and command-r-plus, and 98% for gpt-4o-mini) and Generation X (74% for claude-3.5-sonnet, 48% for gpt-4o-mini, 68% for llama-3.1-70b, and 94% for command-r-plus) individuals. Llama-3.1-70b outputs a higher percentage of heterosexual responses for Generation Alpha (80%) as well. On the other hand, gpt-4o-mini provides a much higher percentage of bisexual responses for Millennials (98%) and Generation Z (98%).
@@
When provided with explicit prompts, the LLMs also output varying occupation distributions based on the gender, age, or ethnic group mentioned in the prompt. The most popular occupation for Baby Boomers, "teacher," is the same across all four models (46% for claude-3.5-sonnet, 90% for gpt-4o-mini, 38% for llama-3.1-70b, and 26% for command-r-plus). Notably, for many of the models, the occupations common among the Generation Alpha texts are unique or have higher proportions compared to the other age groups. For example, popular Generation Alpha occupations include "student" (38.0%) and "digital content creator" (12.0%) for gpt-4o-mini; "student" (68%), "robotics engineer" (10%), and "environmental scientist" (4%) for llama-3.1-70b; and "entreprenuer" (20%), "social media influencer" (16%), and influencer (12.0%) for command-r-plus.
Each model provides different popular occupations for male and female prompts. To illustrate, for gpt-4o-mini, "teacher" (54.0%) and software engineer (24.0%) are the most popular occupations for males, while "environmental scientist" (44.0%) and "graphic designer" (24.0%) are the most popular occupations for females. On the other hand, command-r-plus’s most popular male occupations are "lawyer" (30.0%), "finance" (16.0%), and "financial analyst" (12.0%), while the most popular female occupations are "ceo" (18.0%), "prima ballerina (10.0%), and "executive" (10.0%).
@@ Discussion
The consistent overrepresentation of LGBTQ identities, especially among younger generations, highlights the need to inspect training data and reevaluate model alignment strategies to ensure they reflect real-world distributions more accurately. Additionally, providing greater transparency around how models are aligned would help users better understand the potential biases present in their outputs. While inclusivity is valuable, inflated portrayals unintentionally distort social realities, leading to skewed perceptions or reinforcement of stereotypes.
@@
Refusing to answer inappropriate prompts decreases the risk of an LLM generating harmful or dangerous content. While certain prompts, such as those asking a model to generate violent or offensive texts, are obviously inappropriate, the definition of what makes a prompt appropriate may be less clear in other cases. During the alignment and fine-tuning processes, LLM developers must carefully weigh the balance between effectively serving users’ needs and having guardrails in place to prevent a model from generating harmful content. Future research should investigate the effectiveness of various model alignments and fine-tuning techniques in achieving this balance.
@@ Methods
All models were set with a temperature of 0.7 and top_p of 0.9. The rest of the attributes were default, with the exception of Claude-3.5-Sonnet, where "max_tokens" was set to 1000. In each generation, we asked the model to provide an attributes section containing the attributes of the individual, which we name output attributes. These output attributes contain political affiliation (conservative, liberal, neutral), religion (Christian, Buddhist, Hindu, Jewish, Muslim, unaffiliated), sexual orientation (heterosexual, homosexual, bisexual), and socioeconomic status (lower-class, middle-class, upper-class). Some outputs from the models were not part of these categories, so we manually fit them to such categories (e.g., “lesbian” was changed to “homosexual”, "Catholic" was changed to "Christian"). For sexual orientation, the categories were then grouped into "heterosexual" and "LGBTQ" for ease of analysis with real-world population percentages. The generated attributes were verified to appear in the model’s descriptions through a combination of keyword searches and manual review.
@@
We carefully selected four LLMs for our evaluation and chose these models based on parent company, consumer usage, and LLM’s ability to consistently generate text that followed our instructions and was easy to parse. Our chosen models include claude-3.5-sonnet from Anthropic [37], gpt-4o-mini from OpenAI [38], command-r-plus from Cohere [39], and llama-3.1-70b from Meta [40]. All models come from different companies, are the flagship models in those companies (as of August 2024)[37, 38, 39, 40], and followed our instructions carefully by including an "attributes" section within their output that accurately reflected the generated paragraph.
@@
To evaluate an LLM’s stereotype bias towards a specific demographic group, we compare each of the input attributes’ generated distributions with respect to a given output category. For example, to determine whether an LLM displays socioeconomic stereotype bias against a certain gender, we compare the socioeconomic status distributions of the LLM-generated texts for each gender. This analysis will tell us if a model treats a certain input attribute (e.g., male) differently from its complementary input attribute(s) (e.g., female).
@@
For deviation bias, we compare the distribution of output attributes in the LLM-generated texts for a specific input demographic group with the real-world, ground truth distribution of attributes for that demographic group. This test is crucial because the real-world distribution of a certain demographic attribute (e.g., political affiliation) often depends on our chosen input attributes (e.g., age). The deviation bias metric complements the stereotype bias metric, providing a clearer view of the model’s true biases. To calculate the deviation bias and compare the demographic distributions of the LLM-generated texts with the real-world, ground truth demographic distributions, we perform a binomial test on the frequency of each output attribute. We assess whether the observed frequency for each attribute significantly deviates from what we expect based on the real-world frequency. This method enables us to account for subtle differences that might not be captured by the stereotype bias metric alone, ensuring that the evaluation reflects realistic variations rather than misinterpreting them as bias.
@@
Funding
The authors declare that no funds were received.